### v1

# Determination of the Néel vector in rutile altermagnets through x-ray magnetic circular dichroism: the case of MnF2

Abstract
We present a numerical simulation of the X-ray magnetic circular dichroism (XMCD) at the $L_{2,3}$ edge of Mn in altermagnetic MnF2 using a combination of density functional + exact diagonlization of an atomic model. We explore how the dichroic spectra vary with the light propagation vector and the Néel vector. We show how XMCD in rutile structures can be employed to determine the orientation of the Néel vector. An exact relationship between the XMCD spectra for different orientation of the Néel vector valid in the absence on the valence spin-orbit coupling and core-valence multipole interaction is derived and its approximate validity demonstrated by numerical calculation for the full Hamiltonian.
Introduction. Altermagnets, a new species on the magnetism evolutionary tree, have split from antiferromagnets recently [1, 2]. Owing to their space-group symmetry altermagnets facilitate the presence of spin-polarized bands [3, 4, 1, 5, 6, 7, 8, 2, 9, 10, 11], anomalous Hall effect [12, 5, 13, 14, 15, 9, 16, 17], odd magneto-optical effects [14, 18, 19] and number of other phenomena with odd Néel vector dependence. The key ingredient is the rotation symmetry that connects atoms on different magnetic sublattices, instead of translation or inversion, and makes the time-reversed states with the opposite sign of the Néel vector macroscopically distinct.
Much of the early studies on altermagnetism focused on RuO2 [3, 20, 21, 22, 23], a metal with the rutile structure. Despite considerable theoretical and experimental effort the altermagetism of RuO2 is far from understood and magnetic order in bulk RuO2 remains controversial [24, 25, 26, 27]. Moreover a large magnetic field is needed to tilt the magnetic moments away from the $[001]$ easy-axis direction in order to allow finite odd magneto-optical effects such as circular dichroism [18, 28] or the anomalous Hall effect [20].
These effects are facilitated by the spin-orbit coupling (SOC), which allows reciprocal-space spin textures (spin-polarized bands) of non-relativistic origin to affect the charge response of an altermagnet. However, SOC acts also in the opposite direction and coupling to the orbital degrees of freedom breaks the spin SU(2) symmetry giving rise to the magneto-crystalline anisotropy or a weak ferromagnetism due to canting of the local moments. This disturbs the purely altermagnetic signal and must be taken into account in the analysis of the experimental data [20].
The x-ray circular dichroism (XMCD) has the advantage that SOC is naturally separated into a large core contribution and a small valence contribution. While the latter is responsible for the above ’parasitic’ effects, it has a marginal effect on the XMCD spectra in lighter elements such as $3d$ metals [19, 29]. Given the above uncertainty concerning the magnetism of RuO2 it is desirable to establish the behavior of XMCD in a isostructural material.
MnF2 is a perfect candidate [6]. It Crystallizes in rutile structure and its magnetism is well studied. It orders antiferromangetically around 67 K [30] with magnetic moments are aligned along the $c$-axis [31] similar to RuO2. In contrast to metallic RuO2, MnF2 is a Mott insulator where the Mn2+ configuration gives rise to a large spin moment of $S\approx 5/2$ and an orbital singlet. Together with weak SOC of Mn this results in a negligible single-ion anisotropy. The magneto-crystalline anisotropy with $[001]$ easy axis arises from the long-range dipolar interaction [32, 33]. The spin-flop transition, at which the moments rotate approximately into the $ab$-plane, takes place in the field $\sim 9$ T [33, 34] oriented along the $c$-axis.
[…]
Figure 1: The rutile structure of MnF2 with various orientations of local spin moments.
The XMCD spectrum $\Delta F(\omega)=F^{+}(\omega)-F^{-}(\omega)$ is the difference of absorption of the right and left circularly polarized light propagating along the direction $\hat{\mathbf{k}}$, which is obtained by the Fermi golden rule
$F^{\pm}(\omega;\hat{\mathbf{k}},\mathbf{L})=\sum_{f}\left|\expectationvalue{f_{\mathbf{L}}|\hat{T}^{\pm}_{\hat{\mathbf{k}}}|i_{\mathbf{L}}}\right|^{2}\!\delta\left(\omega\!-\!E_{fi;\mathbf{L}}\right).$ (1)
Here $|i_{\mathbf{L}}\rangle$ and $|f_{\mathbf{L}}\rangle$ are the eigenstates of the Hamiltonian, $E_{fi;\mathbf{L}}$ is the excitation energy, and $\hat{T}^{\pm}_{\hat{\mathbf{k}}}$ are the dipole operators for the right- and left-hand polarization with respect to the propagation vector $\hat{\mathbf{k}}$  [^1] In the following we use the hat symbol $\hat{\mathbf{v}}$ to indicate the direction of vector $\mathbf{v}$, i.e., $\mathbf{v}=a\hat{\mathbf{v}}$ for some $a>0$.. Thanks to the immobility of the core hole the x-ray absorption spectrum is a sum over the site contributions. In the dipole approximation the dependence on $\mathbf{L}$ and $\hat{\mathbf{k}}$ has the form ${\Delta F(\omega;\hat{\mathbf{k}},\mathbf{L})=2\operatorname{Im}\mathbf{h}_{\mathbf{L}}(\omega)\cdot\hat{\mathbf{k}}}$. The axial vector ${\mathbf{h}(\omega)=(\sigma^{a}_{zy}(\omega),\sigma^{a}_{xz}(\omega),\sigma^{a}_{yx}(\omega))}$ representing the antisymmetric part of the conductivity tensor $\boldsymbol{\sigma}(\omega)$ is the finite frequency equivalent of the Hall vector [45, 1, 18].
Figure 2: The XAS calculated for the two circular polarizations (red and blue) at the Mn $L_{2,3}$ edge together with the XMCD intensities (shaded) for different orientations of the Néel vector $\mathbf{L}$ and x-ray propagation vector $\bf k$. The calculated spectral intensities are broadened by Lorentizan of 0.15 eV (HWHM). The experimental Mn $L_{2,3}$-edge XAS spectrum taken from Ref. 40 is shown for comparison.
Figure 3: The orientation of the Hall vector $\mathbf{h}$ (black arrow) and the local spin moments (blue arrows) in the $ab$-plane of the rutile structure.
Figure 4: The Mn $L_{3}$-edge XMCD intensities calculated for various angles $\varphi$, with 0 and 45 corresponding to $\hat{\mathbf{L}}=[110]$ and $[010]$, respectively. The valence SOC and CV interaction is switched off in (a,b). The XMCD intensities in the panels (a,b) collapse on a single curve upon division by $\cos(2\varphi)$ and $sin(2\varphi)$, respectively, as required by (4). (c,d) The same calculations as in (a,b) with full Hamiltonian. (e,f) The spectra from (c,d) divided by $\cos(2\varphi)$ and $sin(2\varphi)$ demonstrate that (4) is approximately fulfilled also with the full Hamiltonian.
Results. The symmetry of the rutile structure [5, 18] implies that XMCD is not allowed for $\mathbf{L}\parallel c$, i.e., $\mathbf{h}=0$. However, it is allowed if $\mathbf{L}$ has a finite projection in the $ab$-plane. First, we consider an idealized situation with magnetic moments entirely in the $ab$-plane, i.e., $\mathbf{L}\perp c$ and $\mathbf{m}_{1}+\mathbf{m}_{2}=0$. Figure 2 shows the XAS and XMCD spectra at the Mn $L_{2,3}$ edges calculated for the Néel vectors $\hat{\mathbf{L}}=[110]$ and $[010]$. The corresponding Hall vectors are parallel ($\hat{\mathbf{h}}=[110]$) and perpendicular ($\hat{\mathbf{h}}=[100]$) to $\mathbf{L}$, respectively [18]. The XAS line shape agrees well with the experimental data [40]. Relatively large XMCD intensities are predicted at both the Mn $L_{2}$ and $L_{3}$ edges in the present theory, Figs. 2ab.
The similarity of the spectra in Figs. 2ab is not accidental. As shown in Ref. 18 the two spectra are related by symmetry if the valence SOC and the multipole part of the core-valence (CV) interaction are neglected. This raises questions about the general orientation of $\mathbf{L}$ in the $ab$-plane. We will address this question by extending the analysis of Ref. 18.
In the absence of the valence SOC and the CV interaction the valence spin is decoupled from the rest of the system and the states (including the excited ones) corresponding to different orientation of $\mathbf{L}$ are related by a valence spin rotation. Following Ref. 18 we fix the coordinate system so that $\hat{\mathbf{x}}=\hat{\mathbf{k}}$ and $\hat{\mathbf{z}}=[001]$ as depicted in Fig. S3 [39], and use dipole operators in the helicity basis along $\hat{\mathbf{z}}$, which is also the quantization axis for spin. Eq. 1 for this coordinates becomes
$\begin{split}\Delta F(\varphi,\alpha)=&\sum_{f}\expectationvalue{f_{\varphi,\alpha}|\hat{T}^{+}-\hat{T}^{-}|i_{\varphi,\alpha}}\expectationvalue{i_{\varphi,\alpha}|\hat{T}^{0}|f_{\varphi,\alpha}}\\ &\times\delta\left(\omega\!-\!E_{fi}\right)+c.c.\\ \equiv&(T^{+}-T^{-})\overline{T^{0}}+c.c.,\end{split}$ (2)
where the angles $\varphi$ and $\alpha$ capture the orientation of the MnF6 octahedra and the local moment with respect to the light propagation vector $\hat{\mathbf{k}}$. The dependence of single site XMCD on $\varphi$ and $\alpha$ has a simple form [18, 39]
$\begin{split}\Delta F(\varphi,\alpha)=&(e^{i\alpha}T_{\uparrow}^{+}-e^{-i(2\varphi-\alpha)}T_{\uparrow}^{-})\overline{T_{\downarrow}^{0}}\\ &+(e^{i(2\varphi-\alpha)}T_{\downarrow}^{+}-e^{-i\alpha}T_{\downarrow}^{-})\overline{T_{\uparrow}^{0}}+c.c.\end{split}$ (3)
Upon summation over the two Mn sites we get
$\begin{split}&\tfrac{1}{2}\left(\Delta F(\varphi,\alpha)+\Delta F(\varphi+\tfrac{\pi}{2},\alpha+\pi)\right)\\ &=e^{i(2\varphi-\alpha)}T_{\downarrow}^{+}\overline{T_{\uparrow}^{0}}-e^{-i(2\varphi-\alpha)}T_{\uparrow}^{-}\overline{T_{\downarrow}^{0}}+c.c.\\ &=A(\omega)\cos(2\varphi-\alpha),\end{split}$ (4)
where the bottom line comes from the fact that (4) must vanish for $\varphi=0$ and $\alpha=\tfrac{\pi}{2}$. In Fig. 4 we check the validity of (4) by explicit calculation in Figs. 4ab.
Eq. 4 has a simple geometrical interpretation. If one starts with $\mathbf{L}$ pointing along $[110]$ or $[1\bar{1}0]$ and rotates it along the $c$-axis, the corresponding Hall vector $\mathbf{h}(\omega)$ rotates by the same angle in the opposite direction, see Fig. 3b, while its $\omega$-dependence remains unchanged, i.e., $\hat{\mathbf{h}}$ is a mirror image of $\hat{\mathbf{L}}$ with respect to $(110)$ or $(1\bar{1}0)$ plane  [^2] The choice of $(110)$ vs $(1\bar{1}0)$ plane fixes the sign of the effect and depends on the assignment of $\mathbf{m}_{1}$ and $\mathbf{m}_{2}$ to the lattice sites.
$\mathbf{h}_{\mathbf{L}}(\omega)=A(\omega)\mathcal{M}_{[110]}\hat{\mathbf{L}}.$ (5)
This formula applies to any structure with the magnetic sublattices related by four-fold rotation axis and another rotation axis perpendicular to it, which determines the zero of $\varphi$ in (4). Without the latter condition a $B(\omega)\sin(2\varphi-\alpha)$ contribution to (4) may appear.
Next, we assess the validity of Eq. 4 in the presence of both the valence SOC and the multipole part of the core-valence interaction. To this end we vary the orientation of $\mathbf{L}$ within the $ab$-plane and compute the XMCD spectra for $\hat{\mathbf{k}}=\hat{\mathbf{L}}$ and $\hat{\mathbf{k}}=\hat{\mathbf{L}}\times\hat{\mathbf{z}}$. This corresponds to fixing $\alpha=0$ and $\alpha=90^{\circ}$, respectively, in the geometry of Fig. 3. The corresponding Mn $L_{3}$-edge XMCD spectra in Figs. 4cd indeed follow the $\cos 2\varphi$ and $\sin 2\varphi$ dependencies. Upon dividing by the trigonometric functions the spectra collapse on almost identical curves shown in Figs. 4ef. It is not clear why the rescaled spectra group into two groups. The accurate adherence of the XMCD spectra to Eq. 5 enables the determination of the Néel vector orientation within the $ab$-plane through the measurement of XMCD’s angular dependence.
Though the easy axis of MnF2 is parallel to [001], the Néel vector $\mathbf{L}$ can be flopped into the $ab$-plane by a field of 9–10 T [34, 33] along the $c$-axis, which results in canting of the Mn moments $\mathbf{S}_{1}$ and $\mathbf{S}_{2}$ along the $c$-axis. The net magnetization along the $c$-axis for fields close to the spin-flop transition was estimated to $\mu_{z}=0.3~{}\mu_{B}$ and 0.5 $\mu_{B}$ in the two Mn sites [33]. In Fig. 5(a), we simulate the effect of canting on the XMCD spectra for $\hat{\mathbf{L}}=[110]$. The net magnetization along the $c$-axis gives rise to a finite $h_{z}$ component of the Hall vector. The magnitude of XMCD for $\hat{\mathbf{k}}=[001]$ in Fig. 5(b) is comparable to the purely altermagnetic effect for $\hat{\mathbf{k}}=[110]$ despite the out-of-plane component of the local magnetic moments being order of magnitude smaller than in-plane one. Similar to the experimental observation on MnTe [19] the altermagnetic ($\hat{\mathbf{k}}=[110]$) and ferromangetic ($\hat{\mathbf{k}}=[001]$) components of the XMCD spectra exhibit distinct shapes.
Finally, we calculate the linear dichroism (XMLD) in Fig. 6, which provides the standard tool to determine the direction of the Néel vector up to a sign. The shape of the XMLD spectra is distinct from the XMCD spectra and thus possible signal contamination due to an imperfect polarization in experimental realization should be easy to eliminate.
Figure 5: The XMCD intensities at the Mn $L_{2,3}$ edge for (left) x-ray propagation vector $\hat{\mathbf{k}}=[110]$ and (right) $\hat{\mathbf{k}}=[001]$ for the Neél vector $\hat{\mathbf{L}}=[110]$ with a small tilt of the Mn magnetic moment along the $z$-axis of $\mu_{z}=0.5~{}\mu_{B}$ at Mn1 (top) and $\mu_{z}=0.5~{}\mu_{B}$ at Mn2 (middle). The total XMCD intensities are shown in the bottom panels.
Figure 6: The XMLD intensities at the Mn $L_{2,3}$ edge for the x-ray propagation vector $\hat{\mathbf{k}}=[001]$ and the Néel vector $\hat{\mathbf{L}}=[110]$ (left) and $\hat{\mathbf{L}}=[010]$ (right). The XMLD is defined for the two x-ray polarization vectors $\boldsymbol{\epsilon}~{}||~{}[110]$ (red) and $\boldsymbol{\epsilon}~{}||~{}[1\overline{1}0]$ (blue). A small tilt of the Mn magnetic moment along the $z$-axis with the values ($\mu_{z}$) indicated in the panels is considered in the simulation. The total XMLD intensities are shown in the bottom panels.
Discussion. We compare XMCD in MnF2 and $\alpha$-MnTe. Both compounds are $S=5/2$ altermagnetic Mn2+ insulators and the presence or absence of the effect (orientation of the XMCD Hall vector $\mathbf{h}$) depends strongly on the orientation of the Néel vector $\mathbf{L}$, but behaves differently when $\mathbf{L}$ is rotated in the $ab$-plane. In $\alpha$-MnTe, $\mathbf{h}$ points along the $c$-axis and changes sign [16, 19] at six nodal points. In MnF2, the XMCD spectrum remains approximately unchanged with $\mathbf{h}$ turning in the $ab$-plane in the opposite sense to the rotation of $\mathbf{L}$. In both compounds the valence SOC has a minor impact on the XMCD spectra. On the other hand, the role of CV multipole interaction is very different. In $\alpha$-MnTe, with $\mathbf{h}\perp\mathbf{L}$ geometry, XMCD completely vanishes if CV interaction (and valence SOC) is absent thanks to presence of the local (3-fold) rotation axis parallel to $\mathbf{h}$ [19]. In MnF2 the $\mathbf{h}\perp\mathbf{L}$ geometry takes place for $\mathbf{L}\parallel[100]$, however, there is no local rotation axis parallel to $\mathbf{h}$ and thus XMCD can exists even in absence of the CV interaction. Therefore the key interaction for appearence of XMCD in $\alpha$-MnTe is a mere minor perturbation in MnF2  [^3] It shifts the peak positions, but does not affect the magnitute of XMCD much.. This is also reflected in the maximal magnitudes of the effect, which is in MnF2 is about five times larger than in $\alpha$-MnTe.
We have calculated the x-ray magnetic circular dichroism (XMCD) in the spin-flop phase of MnF2. The spectral dependence of the in-plane x-ray Hall vector of altermagnetic origin is distinct from the field-induced out-of-plane component. We have found a simple relationship between the light propagation vector maximizing the XMCD ($\mathbf{h}$) and the in-plane Néel vector $\mathbf{L}$, which allows a unique determination of $\mathbf{L}$ from the angular dependence of XMCD and thus opens the possibilities such as domain imaging [48].
Acknowledgements.
We thank Karel Výborný, Andriy Smolyanyuk and Anna Kauch for discussions and critical reading of the manusript. This work was supported by JSPS KAKENHI Grant Numbers 21K13884, 21H01003, 23K03324, 23H03817 (A.H.), and by the project Quantum materials for applications in sustainable technologies (QM4ST), funded as project No. CZ.02.01.01/00/22_008/0004572 by Programme Johannes Amos Commenius, call Excellent Research and by the Ministry of Education, Youth and Sports of the Czech Republic through the e-INFRA CZ (ID:90254)(J.K.).