### v2

# Determination of the Néel vector in rutile altermagnets through x-ray magnetic circular dichroism: the case of MnF2

Abstract
We present a numerical simulation of x-ray magnetic circular dichroism (XMCD) at the $L_{2,3}$ edge of Mn in a representative rutile altermagnet MnF2 using a combination of density functional theory + exact diagonalization of the atomic model. We explore how the dichroic spectra vary with the orientation of the light propagation vector and the Néel vector. An exact relationship between the XMCD spectra for different orientations of the Néel vector, valid in the absence of the valence spin-orbit coupling and core-valence multipole interaction, is derived and its approximate validity for the full Hamiltonian verified by numerical calculation. This relationship allows to determine the in-plane orientation of the Néel vector using the XMCD spectra alone.
Introduction. Altermagnets, a new species on the magnetism evolutionary tree, have split from antiferromagnets recently [1, 2]. Owing to their space-group symmetry altermagnets facilitate the presence of spin-polarized bands [3, 4, 1, 5, 6, 7, 8, 2, 9, 10, 11], anomalous Hall effect [12, 5, 13, 14, 15, 9, 16, 17], odd magneto-optical effects [14, 18, 19] and number of other phenomena [20] with odd Néel vector dependence. The key element is the rotational or mirror symmetry, which links atoms on the distinct magnetic sublattices. This is different from the translational or inversion symmetry seen in conventional antiferromagnets and results in time-reversed states with opposite Néel vectors being macroscopically distinct. The orientation of the Néel vector, including its sign, is thus an important question. Here, we show that the x-ray circular dichroism (XMCD) alone can answer it in some structures.
Much of the early studies on altermagnetism focused on RuO2 [3, 21, 22, 23, 24], a metal with the rutile structure. Despite a considerable theoretical and experimental effort the altermagetism of RuO2 is far from understood as the magnetic order in a bulk RuO2 remains controversial [25, 26, 27, 28]. Moreover a large magnetic field is needed to tilt the magnetic moments away from the $[001]$ easy-axis direction in order to allow finite odd magneto-optical effects such as circular dichroism [18, 29] or the anomalous Hall effect [21].
These effects are facilitated by the spin-orbit coupling (SOC), which allows the spin long-range order of non-relativistic origin to influence the current response, observed in transport and optical experiments. However, SOC acts also in the reverse direction. The coupling of spins to the current breaks the spin SU(2) symmetry, leading to the magneto-crystalline anisotropy or possibly inducing weak ferromagnetism as a result of canting of the local moments [30]. The latter disturbs the fully compensated magnetic state of an altermagnet and must be taken into account in the analysis of the experimental data [21]. XMCD takes advantage of SOC being naturally separated into a dominant core SOC and a minor valence SOC. The latter may give rise to a weak ferromagnetic XMCD signal, but its contribution to the XMCD spectra in lighter elements such as $3d$ metals is marginal [19, 31].
Given the aforementioned uncertainty concerning the magnetism of RuO2 [27] it is desirable to establish the behavior of XMCD in a isostructural material that possesses a well-established magnetic order. MnF2 is a perfect candidate [6]. It crystallizes in the rutile structure and its magnetism is thoroughly studied. The antiferromagnetic order sets in at around 67 K [32] with magnetic moments along the $[001]$ direction [33, 34, 35] similar to RuO2. In contrast to the metallic RuO2, MnF2 is a Mott insulator where the Mn2+ configuration gives rise to a large spin moment of $S\approx 5/2$ and an orbital singlet. Together with a small SOC in the $3d$ shell this results in a weak single-ion anisotropy. When a strong enough magnetic field is applied along the $[001]$ direction, the moments reorient themselves perpendicularly to it, that is, into the $ab$-plane. Meanwhile, the sublattice magnetizations stay antiparallel to each other with a slight tilt towards the field direction. The field required for this spin-flop transition is around 9 T [35, 36]. The direction of the Néel vector within the $ab$-plane is unknown.
[…]
Figure 1: The rutile structure of MnF2 with various orientations of local spin moments.
The XMCD spectrum $\Delta F(\omega)=F^{+}(\omega)-F^{-}(\omega)$ is the difference of the absorption spectra for the right and left circularly polarized light propagating along the direction $\hat{\mathbf{k}}$, obtained by the Fermi golden rule
$F^{\pm}(\omega;\hat{\mathbf{k}},\mathbf{L})=\sum_{f}\left|\expectationvalue{f_{\mathbf{L}}|\hat{T}^{\pm}_{\hat{\mathbf{k}}}|i_{\mathbf{L}}}\right|^{2}\!\delta\left(\omega\!-\!E_{fi;\mathbf{L}}\right).$ (1)
Here $|i_{\mathbf{L}}\rangle$ and $|f_{\mathbf{L}}\rangle$ are the eigenstates of the Hamiltonian, $E_{fi;\mathbf{L}}$ is the excitation energy  [^1] The spectra are calculated relative to the $L_{3}$ edge, the absolute excitation energies are obtained by matching the XAS spectra with experiment., and $\hat{T}^{\pm}_{\hat{\mathbf{k}}}$ are the dipole operators for the right- and left-hand polarization with respect to the propagation vector $\hat{\mathbf{k}}$  [^2] In the following we use the hat symbol $\hat{\mathbf{v}}$ to indicate the direction of vector $\mathbf{v}$, i.e., $\mathbf{v}=a\hat{\mathbf{v}}$ for some $a>0$.. Thanks to the immobility of the core hole the x-ray absorption spectrum is a sum over the site contributions. In the dipole approximation the dependence on $\mathbf{L}$ and $\hat{\mathbf{k}}$ has the form ${\Delta F(\omega;\hat{\mathbf{k}},\mathbf{L})=2\operatorname{Im}\mathbf{h}_{\mathbf{L}}(\omega)\cdot\hat{\mathbf{k}}}$. The axial vector ${\mathbf{h}(\omega)=(\sigma^{a}_{zy}(\omega),\sigma^{a}_{xz}(\omega),\sigma^{a}_{yx}(\omega))}$, representing the antisymmetric part of the conductivity tensor $\boldsymbol{\sigma}(\omega)$, is the finite frequency equivalent of the Hall vector [48, 1, 18].
Figure 2: The XAS calculated for the two circular polarizations (red and blue) at the Mn $L_{2,3}$ edge together with the XMCD intensities (shaded) for different orientations of the Néel vector $\mathbf{L}$ and x-ray propagation vector $\bf k$. The calculated spectral intensities are broadened by Lorentzian of 0.15 eV (HWHM). The experimental Mn $L_{2,3}$-edge XAS spectrum taken from Ref. 42 is shown for comparison (the experimental baseline was offset for the sake of clarity.)
Figure 3: The orientation of the Hall vector $\mathbf{h}$ (black arrow) and the local spin moments (blue arrows) in the $ab$-plane of the rutile structure.
Figure 4: The Mn $L_{3}$-edge XMCD intensities calculated for various angles $\varphi$, with 0 and 45 corresponding to $\hat{\mathbf{L}}=[110]$ and $[010]$, respectively. The valence SOC and CV interaction is switched off in (a,b). The XMCD intensities in panels (a,b) collapse onto a single curve upon division by $\cos(2\varphi)$ and $sin(2\varphi)$, respectively, as required by (4). In (c,d) the same spectra as in (a,b) calculated with full Hamiltonian are shown. Panels (e,f) show the spectra from (c,d) divided by $\cos 2\varphi$ and $\sin 2\varphi$, respectively.
Results. Symmetry of the rutile structure [5, 18] implies that XMCD is not allowed for $\mathbf{L}\parallel c$, i.e., $\mathbf{h}=0$. However, it is allowed if $\mathbf{L}$ has a finite projection in the $ab$-plane. First, we consider an idealized situation with magnetic moments entirely in the $ab$-plane, i.e., $\mathbf{L}\perp c$ and $\mathbf{m}_{1}+\mathbf{m}_{2}=0$. Figure 2 shows the XAS and XMCD spectra at the Mn $L_{2,3}$ edges calculated for the Néel vectors $\hat{\mathbf{L}}=[110]$ and $[010]$. The corresponding Hall vectors are parallel ($\hat{\mathbf{h}}=[110]$) and perpendicular ($\hat{\mathbf{h}}=[100]$) to $\mathbf{L}$, respectively [18]. The spectra for $\hat{\mathbf{h}}=[001]$ a can be found in SM [41]. The XAS line shape agrees well with the experimental data [42]. Relatively large XMCD intensities are predicted at both the Mn $L_{2}$ and $L_{3}$ edges in the present theory, Figs. 2ab.
The similarity of the spectra in Figs. 2ab is not accidental. As shown in Ref. 18 the two spectra are related by symmetry if the valence SOC and the multipole part of the core-valence (CV) interaction are neglected. This raises questions about the general orientation of $\mathbf{L}$ in the $ab$-plane. In the following we extend the analysis of Ref. 18 and show that there is a unique relationship between the in-plane orientation of the Néel vector $\mathbf{L}$ and the XMCD spectrum characterized by the Hall vector $\mathbf{h}_{\mathbf{L}}(\omega)$. By reversing this relationship, the measured $\mathbf{h}_{\mathbf{L}}(\omega)$ can be used to determine an unknown orientation of $\mathbf{L}$ in the sample.
In the absence of the valence SOC and the CV interaction, the valence spin is decoupled from the rest of the system. The states (including excited states) corresponding to different orientations of $\mathbf{L}$ are related by a valence spin rotation. Following Ref. 18 we fix the coordinate system so that $\hat{\mathbf{x}}=\hat{\mathbf{k}}$ and $\hat{\mathbf{z}}=[001]$, see SM [41], and use dipole operators in the helicity basis along $\hat{\mathbf{z}}$, which is also the quantization axis for spin. Eq. 1 for this coordinates becomes   [^3] Below, we do not show the frequency dependence and the summation over $f$ explicitly for sake of readability.
$\begin{split}\Delta F(\varphi,\alpha)=&\sum_{f}\expectationvalue{f_{\varphi,\alpha}|\hat{T}^{+}-\hat{T}^{-}|i_{\varphi,\alpha}}\expectationvalue{i_{\varphi,\alpha}|\hat{T}^{0}|f_{\varphi,\alpha}}\\ &\times\delta\left(\omega\!-\!E_{fi}\right)+c.c.\\ \equiv&(T^{+}-T^{-})\overline{T^{0}}+c.c.,\end{split}$ (2)
where the angles $\varphi$ and $\alpha$ capture the orientation of the MnF6 octahedra and the local moment, respectively, relative to the light propagation vector $\hat{\mathbf{k}}$. The dependence of the single site XMCD on $\varphi$ and $\alpha$ has a simple form [18, 41]
$\begin{split}\Delta F(\varphi,\alpha)=&(e^{i\alpha}T_{\uparrow}^{+}-e^{-i(2\varphi-\alpha)}T_{\uparrow}^{-})\overline{T_{\downarrow}^{0}}\\ &+(e^{i(2\varphi-\alpha)}T_{\downarrow}^{+}-e^{-i\alpha}T_{\downarrow}^{-})\overline{T_{\uparrow}^{0}}+c.c.\end{split}$ (3)
Upon summation over the two Mn sites we get
$\begin{split}&\tfrac{1}{2}\left(\Delta F(\varphi,\alpha)+\Delta F(\varphi+\tfrac{\pi}{2},\alpha+\pi)\right)\\ &=e^{i(2\varphi-\alpha)}T_{\downarrow}^{+}\overline{T_{\uparrow}^{0}}-e^{-i(2\varphi-\alpha)}T_{\uparrow}^{-}\overline{T_{\downarrow}^{0}}+c.c.\\ &=A(\omega)\cos(2\varphi-\alpha),\end{split}$ (4)
where the bottom line comes from the fact that (4) must vanish for $\varphi=0$ and $\alpha=\tfrac{\pi}{2}$. In Fig. 4 we check the validity of (4) by an explicit numerical calculation.
Eq. 4 has a simple geometrical interpretation. If one starts with $\mathbf{L}$ pointing along $[110]$ or $[1\bar{1}0]$ and rotates it along the $c$-axis, the corresponding Hall vector $\mathbf{h}(\omega)$ rotates by the same angle in the opposite direction, see Fig. 3b, while its $\omega$-dependence remains unchanged, i.e., $\hat{\mathbf{h}}$ is a mirror image of $\hat{\mathbf{L}}$ with respect to $(110)$ or $(1\bar{1}0)$ plane  [^4] The choice of $(110)$ vs $(1\bar{1}0)$ plane fixes the sign of the effect and depends on the assignment of $\mathbf{m}_{1}$ and $\mathbf{m}_{2}$ to the lattice sites.
$\mathbf{h}_{\mathbf{L}}(\omega)=A(\omega)\hat{\mathbf{h}}=A(\omega)\mathcal{M}_{[110]}\hat{\mathbf{L}}.$ (5)
Note the geometrical meaning of $\hat{\mathbf{h}}$ as the direction of light propagation for which the XMCD signal is maximal. This formula applies to any structure with the magnetic sublattices related by a four-fold rotation axis and another rotation axis perpendicular to it, which determines the zero of $\varphi$ in (4). Without the latter condition an additional $B(\omega)\sin(2\varphi-\alpha)$ contribution to (4) may appear.
Next, we assess the validity of Eq. 4 in the presence of both the valence SOC and the CV interaction. To this end we vary the orientation of $\mathbf{L}$ within the $ab$-plane and compute the XMCD spectra for $\hat{\mathbf{k}}=\hat{\mathbf{L}}$ and $\hat{\mathbf{k}}=\hat{\mathbf{L}}\times\hat{\mathbf{z}}$. This corresponds to varying $\varphi$ while fixing $\alpha=0$ and $\alpha=90^{\circ}$, respectively, in Eq. 4. Without the valence SOC and the CV interaction the numerical XMCD spectra perfectly follow $\cos 2\varphi$ and $\sin 2\varphi$ dependencies given by (4), see Fig. 4ab. With the valence SOC and the CV interaction turned on, the XMCD spectra in Figs. 4cd still follow the $\cos 2\varphi$ and $\sin 2\varphi$ dependencies to a good accuracy. This is demonstrated in Figs. 4ef by dividing the spectra with $\cos 2\varphi$ and $\sin 2\varphi$, respectively, which leads to a collapse on almost identical curves. It is not clear to us why the rescaled spectra group into two groups, i.e., why the differences between the curves within the panels e and f are smaller than the difference between the panels. We can conclude that, although not exact, Eq. 5 is rather fulfilled even for the full Hamiltonian including the valence SOC and the CV interaction.
Although the easy axis of MnF2 is parallel to [001], the Néel vector $\mathbf{L}$ can be flopped into the $ab$-plane by a field of 9–10 T [36, 35] along the $c$-axis, which causes a small canting of the Mn moments $\mathbf{S}_{1}$ and $\mathbf{S}_{2}$ into the [001] direction. The net magnetization along the $c$-axis for fields close to the spin-flop transition was estimated to $\mu_{z}=0.3~{}\mu_{B}$ and 0.5 $\mu_{B}$ in the two Mn sites [35]. In Fig. 5(a), we simulate the effect of canting on the XMCD spectra for $\hat{\mathbf{L}}=[110]$. The net magnetization along the $c$-axis gives rise to a finite $h_{z}$ component of the Hall vector. The magnitude of XMCD for $\hat{\mathbf{k}}=[001]$ in Fig. 5(b) is comparable to the purely altermagnetic effect for $\hat{\mathbf{k}}=[110]$ despite the out-of-plane component of the local magnetic moments being order of magnitude smaller than in-plane one. Similar to the experimental observation on MnTe [19] the altermagnetic ($\hat{\mathbf{k}}=[110]$) and ferromagnetic ($\hat{\mathbf{k}}=[001]$) components of the XMCD spectra exhibit distinct shapes. This result can be used to estimate the impact of potential misalignment in an experimental set-up.
Finally, we calculate the linear dichroism (XMLD) in Fig. 6, which provides the standard x-ray spectroscopic tool to determine the direction of the Néel vector in antiferromagnets [51, 52], but which cannot distinguish the Néel vectors with opposite orientation, which are of particular interest in altermagents. Our aim is to show the distinct profiles of the XMLD and XMCD spectra, which facilitates the identification and removal of any potential signal contamination arising from imperfect polarization in an experimental implementation.
Figure 5: The XMCD intensities at the Mn $L_{2,3}$ edge for (left) x-ray propagation vector $\hat{\mathbf{k}}=[110]$ and (right) $\hat{\mathbf{k}}=[001]$ for the Neél vector $\hat{\mathbf{L}}=[110]$ with a small tilt of the Mn magnetic moment along the $z$-axis of $\mu_{z}=0.5~{}\mu_{B}$ at Mn1 (top) and $\mu_{z}=0.5~{}\mu_{B}$ at Mn2 (middle). The total XMCD intensities are shown in the bottom panels.
Figure 6: The XMLD intensities at the Mn $L_{2,3}$ edge for the x-ray propagation vector $\hat{\mathbf{k}}=[001]$ and the Néel vector $\hat{\mathbf{L}}=[110]$ (left) and $\hat{\mathbf{L}}=[010]$ (right). The XMLD is defined for the two x-ray polarization vectors $\boldsymbol{\epsilon}~{}||~{}[110]$ (red) and $\boldsymbol{\epsilon}~{}||~{}[1\overline{1}0]$ (blue). A small tilt of the Mn magnetic moment along the $z$-axis with the values ($\mu_{z}$) indicated in the panels is considered in the simulation. The total XMLD intensities are shown in the bottom panels.
Discussion. Next, we compare XMCD in the spin-flopped phase of MnF2 and $\alpha$-MnTe. Both compounds are $S=5/2$ altermagnetic Mn2+ insulators and the presence or absence of the effect as well as the orientation of the XMCD Hall vector $\mathbf{h}$) depends on the orientation of the Néel vector $\mathbf{L}$, but behaves differently when $\mathbf{L}$ is rotated in the $ab$-plane. In $\alpha$-MnTe, $\mathbf{h}$ points along the $c$-axis and changes sign, vanishes for all frequencies, [16, 19] at six nodal points. In MnF2, the shape of XMCD spectrum remains approximately unchanged, but $\mathbf{h}$ rotates in the $ab$-plane in the opposite sense to the rotation of $\mathbf{L}$. In both compounds the valence SOC has a minor impact on the XMCD spectra. On the other hand, the role of CV multipole interaction is very different due to the different symmetries of the crystal fields in the two compounds.
In $\alpha$-MnTe, with $\mathbf{h}\perp\mathbf{L}$ geometry, XMCD completely vanishes if the valence SOC and the CV interaction are absent. This is caused by the presence of local (3-fold) rotation axis parallel to $\mathbf{h}$ as explained in Ref. 19. The $\mathbf{h}\perp\mathbf{L}$ geometry takes place also in MnF2 for $\mathbf{L}\parallel[100]$, however, there is no local rotation axis parallel to $\mathbf{h}$ and there XMCD is allowed even if the valence SOC and the CV interaction are absent. These terms modify the shape of the spectra, Fig. 4, but do not change the magnitude of XMCD substantially. Therefore the key interaction for the appearance of XMCD in $\alpha$-MnTe is a mere perturbation in MnF2. Different origin of the XMCD is also reflected in the XMCD magnitudes, which in MnF2 is about five times larger than in $\alpha$-MnTe.
We have studied the x-ray magnetic circular dichroism in the spin-flop phase of MnF2. Using an approximate symmetry, we have found a simple relationship between the light propagation vector maximizing the XMCD, $\hat{\mathbf{h}}$, and the in-plane Néel vector $\mathbf{L}$ in the rutile structure, which allows a unique determination of $\mathbf{L}$ from the angular dependence of XMCD. Comparing MnF2 and $\alpha$-MnTe, we have shown that even in isoelectronic compounds XMCD may originate in different terms in the Hamiltonian depending on their symmetries.
Acknowledgements.
We thank Karel Výborný, Andriy Smolyanyuk and Anna Kauch for discussions and critical reading of the manuscript. This work was supported by JSPS KAKENHI Grant Numbers 21K13884, 21H01003, 23K03324, 23H03817 (A.H.), and by the project Quantum materials for applications in sustainable technologies (QM4ST), funded as project No. CZ.02.01.01/00/22_008/0004572 by Programme Johannes Amos Commenius, call Excellent Research and by the Ministry of Education, Youth and Sports of the Czech Republic through the e-INFRA CZ (ID:90254)(J.K.).